VT001

PI: Edwards
Title: VSOPTEST-1

Session: February 1996 (session 1)
Frequency: 18 cm
Mode: b (double-speed mode B)
Obs time (UT): 44 EU 06:00(13/02)-07:30(13/02)
Stations: NshJW
Mode: b (double-speed mode B)
Obs time (UT): 44 EU 13:00(13/02)-15:00(13/02)
Stations: Sh
Processor: Socc
Mode: b (double-speed mode B)
Obs time (UT): 44 11:00(13/02)-13:00(13/02)
Stations: SL
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Metsahovi (Ari Mujunen, Kaj Wiik at Wed Feb 14 05:38:01 EST 1996) Minor failures
Metsähovi participated the first VSOP test to check technical readiness and tape+file logistics only. We DO NOT have a 18cm receiver, so the tape produced contains 22GHz data plus the 1MHz phase cal signal. No correlation proper, I presume, but perhaps at least some kind of data quality check? We repeat: the tape DOES NOT contain 18cm data, 1.3cm instead. Re experiment: all went apparently well. The schedule requested 3c273 until our antenna elevation was about 1 degree (!), though. The log file has been deposited in: VLBEER.bo.cnr.it: USER3:[evn.feb96]vt001log.mh We will put the log file into an ISAS ftp site as soos as we get access information to it. NRAO aspen.aoc.nrao.edu will also follow as soon as we find its user name/password... :-)

Effelsberg (D Graham at Wed Feb 14 08:54:11 EST 1996) Minor failures
Telescope is currently being renovated, spun a tape on VLBA terminal. IF from noise source connected 06-09 UT, after that no input. NO TELESCOPE SIGNAL AT ANY TIME !. No continuous cal, no pcal. VLBA style log sent to vlbaops, tape off today (14th).

Onsala (John Conway at Mon Feb 19 08:32:57 EST 1996) Minor failures
Onsala was scheduled to take part but did not. Emergency repairs were being made to tape drive capstan motor and capstan itself, these were thought to be the source of problems in tape path stability (and indeed seem to the origin of our problems). Repairs were also being made to our antenna at same time.

Simeiz (Alexandre Volvatch at Wed Feb 21 08:26:33 EST 1996) Minor failures
4 passes (06:16-06:40) were omited because source was behind the hill. 9 passes (17:42-12:51) were omited because we had not power.

Urumqi (Lee Suqin at Wed Feb 21 09:10:36 EST 1996) Success
Cal temprature is 45k and the system temperature is very high because there is one contactor in Lo being loosn. It is sown. Observation is OK.

Jodrell (Dave Henstock at Thu Feb 22 06:40:05 EST 1996) Success
Apparently no problems in either part except late start (~2 minutes) of second part. Tsys is in Jy, need to multiply measured Tsys by 1.23 for better calibration.

Medicina (Venturi Tiziana at Thu Feb 22 09:57:23 EST 1996) Success
First VSOP test carried out successfulyy in Medicina.

Noto (Corrado Trigilio at Sat Feb 24 06:24:33 EST 1996) Success
First observation in VLBA mode in Noto. Used for the first time FS-9.1.0. Apparentelly all the observations carried out successfully.

Westerbork (foley at Mon Feb 26 11:44:19 EST 1996) Success
Only observed 0600-0730 UT. Looks OK locally

Shanghai (Huang Xinyong at Fri Mar 1 04:35:59 EST 1996) Failure
Observations did not take place, due to problems with upgrading F.S from 8.2 to 9.1.